### v1

# Task and Motion Planning in Hierarchical 3D Scene Graphs

Abstract
Recent work in the construction of 3D scene graphs has enabled mobile robots to build large-scale hybrid metric-semantic hierarchical representations of the world. These detailed models contain information that is useful for planning, however how to derive a planning domain from a 3D scene graph that enables efficient computation of executable plans is an open question. In this work, we present a novel approach for defining and solving Task and Motion Planning problems in large-scale environments using hierarchical 3D scene graphs. We identify a method for building sparse problem domains which enable scaling to large scenes, and propose a technique for incrementally adding objects to that domain during planning time to avoid wasting computation on irrelevant elements of the scene graph. We test our approach in two hand crafted domains as well as two scene graphs built from perception, including one constructed from the KITTI dataset. A video supplement is available at https://youtu.be/63xuCCaN0I4.
I INTRODUCTION
We aim to enable an autonomous agent to solve large-scale Task and Motion Planning (TAMP) problems in real-world environments. In order to solve a TAMP problem, an abstract planning domain is needed, which efficiently and accurately represents the robot’s environment and the available actions. Over the past several years, significant progress has been made in the area of generating hybrid metric-semantic representations of the world using 3D scene graphs. One such example is the Hydra scene graph [12], which jointly represents low-level information such as scene geometry and robot trajectories, and higher level information such as semantically-annotated object instances, places, rooms, and buildings. This environmental abstraction lends itself well to large-scale TAMP problems, as it is capable of storing both higher level abstractions like objects and connectivity of regions needed for task planning, as well as the low-level metric information required to check kinematic feasibility of different actions [12]. However, how best to construct a planning domain from the 3D scene graph representation remains an open question.
The trend in scene graph development has been toward encoding an increasing amount of information in the hierarchical representations. Such detailed information may be crucial to the success of a planner, as the more faithfully any planning domain (regardless of the planning system or formalism) represents the real world, the more likely a found solution will actually be executable by the robot. Moreover, increased representational capacity of the planning domain leads to a wider range of robot goals that can be specified, thus increasing flexibility and generality.
However, as a planning problem instance grows in the number of objects, so too does the complexity of finding a plan. TAMP is PSPACE-Hard [22], so problems can become computationally intractable very quickly as the sizes of the state and action spaces grow [7]. To avoid creating intractable planning problems when converting a 3D scene graph into a planning problem, it is critical to determine which elements of the environment (i.e., which elements of the scene graph) are relevant. Consider, for example, a robot responding to a Chemical, Biological, Radiological, Nuclear, and Explosive (CBRNE) scenario, receiving instructions to inspect and potentially neutralize dangerous objects scattered in a large area, represented as a pre-built scene graph. The robot is able to pass near an object only after it has neutralised and cleared it, and may be instructed to avoid particular regions entirely, to solve different tasks. Depending on the geometry of the scene and the specified goal, only a subset of these potentially dangerous obstacles and regions may ultimately be relevant to finding a plan. But, for a robot building a scene graph representation from perception, it is not at all obvious which elements of the scene graph should be added to a planning domain to ensure a valid plan can be found and executed. If we were to add every potentially relevant element to our planning domain, we would be guaranteed to eventually find a plan which takes our robot to its goal, but at the expense of substantial computational complexity. On the other hand, if we are too aggressive in pruning away seemingly irrelevant elements, we may restrict ourselves to sub-optimal solutions, or fail to find a plan at all.
Previous approaches to the problem of inferring a task-relevant planning domain have relied on representations of connectivity in the scene graph to prune superfluous elements [1]. However, these efforts have been limited to very specific task planning problems, as the pruning approaches employed often remove information necessary for checking the geometric feasibility of plans, or implicitly limit the types of goals that can be specified. Alternative approaches for reducing the planning domain size involve attempting to learn the relevance of planning objects, then incrementally adding objects to the domain according to the learned relevance score until the problem is solvable [18]. Unfortunately, this approach requires training on numerous similar planning problems, and is difficult to generalize to tasks at large scales.
In this work, we propose a novel approach to both enable and accelerate TAMP in large environments (Fig. 1). The first contribution of this work is the definition of a set of properties that a TAMP domain must possess to ensure that any plans returned by solving a problem actually meet the user’s specifications. We then define a method for translating a Hydra scene graph [12] into a planning domain that meets these criteria. Our second contribution is the formulation of a sufficient condition for removing symbols from a planning problem while maintaining feasibility. This condition shows that, for example, many of the places in a complete scene graph can be ignored when formulating certain kinds of planning problems, greatly reducing planning times. This improvement relies on the definition of a hierarchical planner, which ignores certain elements of the scene at the highest level to accelerate discrete planning, while planning over the entire scene at lower levels to ensure constraints on the plan are not violated. The third contribution is a method to further accelerate planning by incrementally identifying objects in the scene as relevant during search according to how they affect the feasibility of certain abstract plans. In contrast to the places component of the domain, finding a small set of objects ahead of time to include in the planning instance while guaranteeing that the problem remains feasible is quite difficult. Instead, each time we fail to solve a sub-problem in an abstract plan, we use the scene graph to identify which objects were responsible. We then augment our domain with those objects, and continue planning. We show the effectiveness of our approach across two hand-crafted domains, in addition to two scene graphs built from real perception.
II Task and Motion Planning in 3D Scene Graphs
Figure 1: An illustration of how we derive our planning representation from a 3D scene graph generated from the KITTI dataset. In (A), we present three levels of the scene graph hierarchy, with the low-level Mesh layer on the bottom, the Places layer above that, and the Objects layer on the top. In (B), we see an isometric view of this graph, giving an insight to the scale, while also plotting the higher level Regions layer. In (C) we show a very simplified version of this scene, where the agent is given a task specified in logic. The agent is tasked with either visiting Place 6 while avoiding Place 1, or visiting Place 5. Given the geometry of the scene, the first clause in this disjunction is un-satisfiable. We also see that Place 5 is partially obstructed by a suspicious object, so the agent must consider either avoiding it (green trajectory), or inspect and neutralize the object (blue trajectory) before continuing to its goal. To solve problems like this one at large scales, we take advantage of the structure of the scene graph to accelerate planning.
TAMP jointly considers elements of high-level task planning [8, 13] and low-level motion planning [15] in an attempt to solve hybrid discrete/continuous, multi-modal planning problems [7].
II-A Task and Motion Planning Preliminaries
A common formalism for encoding planning problems is Planning Domain Definition Language (PDDL). In a PDDL problem, a state $\mathcal{I}$ is a set of facts: instances of boolean functions called predicates $p(\bar{x})\in\mathcal{P}$, which are parameterized by tuples of objects $\bar{x}=[x_{1},...,x_{k}]$, $x\in\mathcal{O}$. Transitions between states are defined by actions $a(\bar{x})\in\mathcal{A}$ (also parameterized by objects) which are expressed as two sets of predicates: preconditions $\text{Pre}(a_{i})$ and effects $\text{Eff}(a_{i})$. An action’s preconditions determine if an instance can be applied from a particular state, and its effects define the set of facts that are added ($\text{Eff}^{+}(a_{i}))$ or removed $(\text{Eff}^{-}(a_{i}))$ from the state. A planning domain is composed of lifted sets of predicates $\mathcal{P}$ and actions $\mathcal{A}$, and a problem instance $P=(\mathcal{P},\mathcal{A},\mathcal{O},\mathcal{I}_{0},\mathcal{G})$ combines a domain with an initial state $\mathcal{I}_{0}$ and a set of goal states $\mathcal{G}$, parameterized by objects $\mathcal{O}$. We will refer to objects in a PDDL problem as symbols when they may be confused with physical objects from a scene graph.
One approach for representing a TAMP problem —which we use in this work— is an extension of PDDL called PDDLStream [6]. A PDDLStream problem instance $(\mathcal{P},\mathcal{A},\mathcal{S},\mathcal{O},\mathcal{I}_{0},\mathcal{G})$ represents the discrete search portion of a TAMP problem in PDDL, with the same representation of predicates, actions, objects, initial state, and set of goal states. In a TAMP problem however, certain actions may have symbol parameterisations that are either cumbersome or intractable to add to a problem instance. For example, consider a move action, which defines robot motion from one configuration to another, and which may have a precondition that requires the existence of a collision-free trajectory between those configurations. Adding all such trajectory symbols between all possible points in the configuration space to the planning domain is impossible, so instead, we define sub-problems called streams $s\in\mathcal{S}$ which can be solved to return symbols that (if a solution exists) can be used to certify the preconditions of the associated action. In the case of a trajectory, the relevant sub-problem/stream might be a motion planner, which, if solved, returns a symbol representing a collision-free trajectory between two configurations that the robot can follow. We refer the reader to [6] for a detailed description of PDDLstream.
Solutions to PDDLStream problems take the form of a sequence of parameterized action instances $\pi=[a_{1}(\bar{x}_{1}),a_{2}(\bar{x}_{2}),...,a_{n}(\bar{x}_{n})]$ that define a plan, where assigned parameters $\bar{x}$ satisfy each action’s constraints [7]. From the parameters of an action, we can derive a motion sequence, which specifies how a robot executes a particular action. For example, from an action plan composed of a sequence of move actions, the corresponding motion sequence would be composed of the trajectories that were solved for by the motion-planning sub-problem and which are parameters of those move actions. Executing that sequence would involve multiple calls to a trajectory controller. We say that two motion sequences are equivalent if they result in the agent acting identically. For any action sequence, there is a corresponding sequence of states $\mathcal{I}_{\pi}=[\mathcal{I}_{0},\mathcal{I}_{1},\mathcal{I}_{2},...,\mathcal{I}_{n}]$, leading from the initial state to a goal state that can be constructed from each action’s effects. For an action plan $\pi$, its corresponding state plan $\mathcal{I}_{\pi}$ is valid if $\mathcal{I}_{i}\in\text{Pre}(a_{i+1})$ for $i=0,...,N-1$, and $\mathcal{I}_{N}\in\mathcal{G}$. Tasks specified in PDDL can be solved by a range of solvers [9, 11], and any state plan found by such a solver is valid by construction. A feasible planning problem is one for which there is a valid solution.
II-B Building 3D Scene Graphs from Perception
In most TAMP settings, the planning domain and problem instance are assumed to be given ahead of time. However, a robot operating in the real world must be able to build its own world model from perception, especially the symbols and predicates that describe the world states. Recently, 3D scene graphs such as those implemented by Hydra [12] or S-Graphs [3] have emerged as promising hierarchical representations that can be built in real-time onboard robots, to provide abstractions that are useful for encoding planning domains.
Our definition of a 3D scene graph, directly based on Hydra  [12], consists of several layers of increasing abstraction (see Fig. 1). Each layer consists of a collection of nodes representing the location and other attributes of an element, with edges connecting nodes within the same layer representing relative spatial constraints and edges between different layers representing an inclusion relationship. The lowest layer of the hierarchy is a semantically-annotated mesh of the scene geometry. The next layer contains objects identified by a semantic image segmentation. The places layer represents navigable regions of the environment based on semantic and geometric properties of the mesh. Places are clustered into groups based on geometric and semantic information, and these groups become nodes in the higher-level regions layer (e.g., rooms in an indoor environment).
Previous work on 3D scene graphs has mainly focused on indoor uses. These representations rely on the Generalized Voronoi Diagram (GVD) [16] to generate places, an abstraction of 3D spatial connectivity, which are not well suited for ground robot navigation. We use an alternate formulation of 2D places in our navigable scene graph. Each of these places represents a 2D polygon with consistent terrain classification, and is connected to adjacent places in the scene graph. As the resolution of the places is much coarser than the mesh resolution, planning over sequences of places can be much faster than planning over the mesh itself, while still retaining important geometric information. The learned metric-semantic method for extracting regions from places presented in Strader et al. [21] is more appropriate for these 2D places compared to the purely geometric method room-based region extraction originally presented in Hughes et al. [12]. Hydra and its region-based clustering extension [21] can construct this map representation in realtime from RGBD sensor data while accounting for odometry drift, enabling large scale, consistent, and information-rich maps.
II-C Inferring the Planning Domain from Scene Graphs
We introduce a method for deriving a planning problem instance from a Hydra scene graph to demonstrate the salient aspects of solving planning problems based on large-scale environments. In general, we generate six classes of predicates that may be relevant for scene graph related planning. 1) Type information derived from the nodes of the graph, where each node is a unary predicate: (Configuration ?c), (Place ?p), and (Object ?o), etc. 2) Agent/object predicates that define the state of the robot and objects: (AtConfig ?c), (AtPlace ?p), (AtRoom ?r), etc. 3) Connection predicates defined by edges in the same level of the graph: (Connected ?n1 ?n2). 4) Hierarchical predicates indicating edges connecting nodes of different levels of the graph: (PoseInPlace ?c ?p), (PlaceInRoom ?p ?r), etc. 5) Preconditions of actions that are certified by solving a stream’s associated sub-problem. For the example move action, the predicate certified by the motion-planner would contain the two configurations, and the trajectory between them: (Trajectory ?c1 ?t ?c2). 6) Additional, problem specific predicates defined by the user that can be used to specify goal states and problem constraints such as which places to visit or which objects to collect.
As a running example for the remainder of this paper, we define an example problem using these predicates, motivated by CBRNE scenarios. In this “Inspection Domain”, an agent can be commanded to visit or avoid certain places, and inspect and neutralise objects that have been marked as suspicious. These suspicious objects cannot be passed until they have been inspected and neutralised. We therefore define problem specific predicates: (VisitedPlace ?p) which indicate the current and past places the robot has been to, and (Safe ?o) or (Suspicious ?o) which describe an object. We define sub-problems for planning motion between two poses, sampling poses for inspecting and neutralizing objects, and sampling poses in a specific place. Goal specifications in this domain can include positive or negated facts based on these predicates. The agent’s available actions are to move between poses in connected places, and to inspect objects from appropriate poses (for simplicity we do not separate the inspect and neutralise actions). Any valid plan is composed of these actions, which at execution time are converted into a motion sequence of FollowPath($t_{i}$) and InspectObject($o_{i}$) primitives for paths $t_{i}$ and objects $o_{i}$. Places and objects are the most numerous nodes in the scene graph, so we need to ensure solving tasks in our domain remains tractable as the number of places and objects grows. The complexity of dealing with places and objects will have the largest implication for the move action, which we present here:
  (:action move
    :parameters (?p1 ?p2 ?c1 ?c2 ?t)
    :precondition (and
              (Trajectory ?c1 ?t ?c2)
              (PoseInPlace ?c1 ?p1)
              (PoseInPlace ?c2 ?p2)
              (Connected ?p1 ?p2)
              (AtPose ?c1)
    :effect (and (AtPose ?c2)
                 (not (AtPose ?c1))
                 (VisitedPose ?c2))))
  (:derived (AtPlace ?p)
    (exists (?pose)
            (and (AtPose ?pose)
                 (PoseInPlace ?pose ?p))))
This action allows the agent to move between its current place $p_{1}$ and any place $p_{2}$ that is connected to $p_{1}$ in the scene graph, and updates the state to note the agent’s current pose and that $p_{2}$ has been visited. We refer to this problem representation as the direct encoding of the Inspection domain.
The direct encoding of the problem has one major drawback: moving between places that are not physically near each other necessitates chaining together potentially many move actions, which requires motion planning in many short segments between neighboring places. For simple navigation, planning in short segments can actually aid in the efficiency of long-distance motion planning, as we will discuss in Sec.III-C. However, relying on a general task planner to construct long-horizon action sequences, any segment of which may be invalidated by the low-level geometry, leads to very inefficient planning in practice. In the Inspection domain for example, solving a motion-planning sub-problem for a trajectory requires accounting for Suspicious objects that may block the way. However, there is no a priori logical connection in the PDDL problem between a trajectory and the objects which might interfere with it. Therefore, choosing the right object to inspect in order to unblock a path is difficult for the task planner. The longer the sequence of move actions needed to reach a goal state, the more potential plans the solver must consider. This scales poorly both in the size of the environment, and as we consider increasingly complex goal specifications (Sec. IV).
III Scalable Scene Graph Planning
Our objective is to both enable, and reduce the computational burden of solving TAMP problems in large environments. We emphasize three main criteria in our development of a DSG planning framework:
  1. 1.
    Generating plans that we know can be executed.
  2. 2.
    Preserving the set of goals that can be specified and solved.
  3. 3.
    Planning at large scale.
Solving problem instances with the direct encoding described above is computationally intractable for large scenes. To address this potential intractability, we first define the characteristics of a planning domain that, if met, would enable sparsification while still meeting the criteria listed above. Next, we propose a method for constructing a simplified planning domain that has these characteristics and enables planning in large scenes. Critically, our method depends on factoring the task and motion planning problem into separate task, navigation, and motion planning problems, aligned with different layers of the scene graph hierarchy. Finally, we describe our approach for planning, using feedback from failed attempts to solve sub-problems to add facts to our domain.
III-A Simplifying Problem Instances
One method for reducing the complexity of a planning problem is to shorten the depth and branching factor of the search that produces a plan. In the case of the direct encoding of our Inspect Domain, one way to do this would be to prune Place symbols from the planning domain that are not relevant to the problem instance, and redefine the move action to allow for travel over longer distances. This pruning would remove possible regions the robot might travel and potentially reduce the depth of search required to find a plan, thus simplifying the search problem. However, identifying symbols that do not impact the solution is in general as hard as solving the planning problem itself, and simply removing places from a problem instance might render the problem as encoded in Section II-C infeasible if it is done naively. In this section, we will show that slight modifications to the domain and motion planning sub-problem lead to straightforward identification of symbols that can safely be removed from the planning problem. We begin by defining a set of symbols that are redundant for a particular goal specification.
Definition 1 (Redundant Symbol).
For a set of domain actions $\mathcal{A}$ and specific goal $\mathcal{G}$, a symbol x is redundant if both of the following hold:
  1. 1.
    For every valid plan $\pi$ where x parameterizes an action, there is another valid plan $\pi^{\prime}$ with equivalent motion sequence, where x is not an action parameter,
  2. 2.
    No action precondition or goal, expressed in negative normal form, contains a universal quantifier that can be parameterized by x.
The brief intuition behind this notion of redundancy is that 1) if any plan involving the symbol yields a motion sequence that can be rewritten without the symbol, the symbol is redundant, and 2) if we solve a planning instance where a redundant symbol has been removed, we would like to know that the plan is still valid in the original problem. The requirements for a symbol to be redundant are quite strong (every plan that uses a symbol must have an alternate plan that does not use the symbol and still results in the same motion sequence), but we will later show that a slight modification to the movement action lets many places have this property. We now prove that we do not lose anything by removing a redundant symbol.
Proposition 1 (Removing Redundant Symbol Preserves Feasibility).
Consider a feasible planning instance $(\mathcal{P},\mathcal{A},\mathcal{S},\mathcal{O},\mathcal{I}_{0},\mathcal{G})$. For a redundant symbol $x\in\mathcal{O}$, we define a related instance $R^{\prime}=(\mathcal{P},\mathcal{A},\mathcal{S},\mathcal{O}^{\prime},\mathcal{I}_{0}^{\prime},\mathcal{G}^{\prime})$ where $x$ has been removed, i.e., $\mathcal{O}^{\prime}=\mathcal{O}\setminus x$ and $\mathcal{I}_{0}^{\prime}$ contains all facts in $\mathcal{I}_{0}$ except those parameterized by $x$, and similarly for $\mathcal{G}^{\prime}$. Let $\Pi_{R}$ denote the set of valid plans for $R$. Then, $\Pi_{R^{\prime}}\subseteq\Pi_{R}$ and $\Pi_{R^{\prime}}\neq\emptyset$.
Proof.
Consider $\pi\in\Pi_{R}$. If $\pi$ does not contain any actions parameterized by $x$, then the same plan $\pi$ is also a valid solution for $R^{\prime}$. Consider the alternative case, where $\pi$ does contain an action parameterized by $x$. By Definition 1, there is another plan $\pi^{\prime}$ with equivalent motion sequence not parameterized by $x$, which is a valid solution for $R^{\prime}$.
Now that we have shown that $\Pi_{R^{\prime}}$ is not empty, we need to show that any valid plan for $R^{\prime}$ is valid for $R$. Consider plan $\pi=[a_{1},...,a_{N}]\in\Pi_{R^{\prime}}$ with corresponding state plan $\mathcal{I}_{\pi}=[\mathcal{I}_{0},...,\mathcal{I}_{N}]$. If the addition of facts $\mathcal{F}$ parameterized by $x$ make $\pi$ invalid, then there must exist a state $\mathcal{I}_{k}$ such that $\mathcal{I}_{k}\cup\mathcal{F}\notin\text{Pre}(a_{k+1})$, which means that $a_{k+1}$ is parameterized by a symbol that did not exist in $R^{\prime}$. Only a universal or existential quantifier in $\text{Pre}(a_{k+1})$ can cause $a_{k+1}$ to be parameterized by an additional symbol. Adding additional facts cannot turn an existentially quantified formula from true to false. By Definition 1, $a_{k+1}$ does not have any universal quantifiers that can be parameterized by $x$ in its precondition. Thus $\pi$ must be valid for $R$.
In the original direct domain encoding presented in Sec. II-C, there are no redundant places, since removing any place implies that no other plans could result in a trajectory going through that place, which may be necessary depending on the geometry of the scene. However, let us consider a domain with a more general move action: moveRelaxed. This action takes $N$ places as parameters and has a precondition that a sub-problem is solved for a trajectory that goes through each of these $N$ places (and possibly others not included in the set of $N$). Trajectory represents this trajectory.
  (:action moveRelaxed
    :parameters (?p1 ... ?pN ?c1 ?t ?c2)
    :precondition (and
        (Trajectory ?p1 ... ?pN ?c1 ?c2 ?t)
        (AtPose ?c1)
    :effect (and (AtPose ?c2)
                 (not (AtPose ?c1))
                 (VisitedPlace ?p1)
                 (VisitedPlace ?pN)
                 (VisitedPose ?c2))))
With these modifications, potentially many places become redundant. For any place that appears in a plan incidentally (i.e., not part of the goal), there now exists an equivalent plan where these intermediate places do not appear.
For example, imagine our robot begins in Place A, and is tasked with (among other things) inspecting an object in place C. A motion sequence corresponding to a plan to move from Place A to Place B, then from Place B to Place C can be equivalent to a sequence generated from a plan to move from Place A to C directly, so it seems that we can remove Place B from the problem. Extrapolating this pruning approach to larger scenes and more complex goals has the potential to vastly reduce planning horizons. However, now consider that the robot was also instructed to avoid Place B. The ability to include a constraint on the goal states of the form (not (VisitedPlace B)) complicates the pruning process. Executing moveRelaxed from Place A to Place C may involve following a trajectory that takes the robot through Place B, even if the goal specifies that Place B should not be visited, whether or not Place B is in problem domain. Technically this is still a valid solution to the planning problem since place B never appears as a parameter to the moveRelaxed action (and therefore (VisitedPlace B) is not an effect), but clearly the domain with a relaxed movement action does not fully capture how the user wants the robot to interact with the environment.
III-B Execution-Consistency
To formalize the discrepancy between what happens when the robot executes a motion sequence and the constraints that we expect a planning problem to impose, we introduce the concept of a verifier function. A verifier function maps motion sub-sequences to sets of PDDL domain facts, and “verifies” which additional domain facts must be implicitly true as a result of the agent executing a motion sequence, even if actually adding these facts to the problem instance during the solving process is undesirable computationally. For example, if we want to consider if an agent’s trajectory avoids certain places, we can imagine a verifier that would return (VisitedPlace $p_{i}$) for each place $p_{i}$ that the continuous trajectory overlaps with. Given a verifier $V$, the facts that hold at each step when executing a motion sequence may be different than expected in the original plan. We denote the facts that would be added by such a verifier applied to the motion sequence associated with $a_{i}$ as $V(a_{i})$, and term this sequence of expanded states the $V$-extended state plan.
Definition 2 (V-Extended State Plan).
For an action plan $\pi=[a_{1},...,a_{n}]$, its corresponding state plan $\mathcal{I}_{\pi}=[\mathcal{I}_{0},...,\mathcal{I}_{n}]$, and hypothetical verifier function $V$, the V-extended state plan for state $\mathcal{I}_{k}$ is:
$\mathcal{I}^{\prime}_{1}=\mathcal{I}_{1}\cup V(a_{1})$
[…]
An extended state $\mathcal{I}^{\prime}_{k}$ is composed of the facts $\mathcal{I}_{k}$ in the initial plan, plus any extra facts that were present in the previous extended state $\mathcal{I}^{\prime}_{k-1}$ other than those removed by action $a_{k}$, plus any facts that would be returned by a verifier applied to action $a_{k}$. Informally, $\mathcal{I}^{\prime}_{k}$ is the state at step $k$ as experienced by the verifier. As discussed in Sec. II-A, any state plan found by a search algorithm is valid by construction. However, a state plan that is augmented with the extra facts that would be produced by a verifier might not be, for example, if the goal specification has a condition that the robot should avoid certain places in the domain.
To address this problem, let us imagine a verifier $V_{place}$ that takes a motion sub-sequence $\mu$, and returns a VisitedPlace fact for each place that intersects with the agent’s position while executing $\mu$. For a place $p$ and a trajectory $t$ to be followed by the motion primitive $\texttt{FollowPath}(t)$, we denote $p\cap t$ the section of $t$ that intersects with $p$. We can then define the verifier as
$\begin{split}V_{place}(\mu)=\{&\texttt{(VisitedPlace p)}~{}|~{}\\ &p\cap t_{i}\neq\emptyset\text{ for }\texttt{FollowPath}(t_{i})\in\mu\}.\end{split}$ (1)
If the motion sequence associated with the action plan would result in the agent visiting a place that we do not expect, then the $V_{place}$-extended state plan would include a VisitedPlace fact that may conflict with the goal. From this idea, we define the concept of execution-consistency, which requires that solutions to the planning problem are still valid after considering the facts from a verifier.
Definition 3 (Execution-consistent).
A domain is execution-consistent with respect to verifier $V$ if, for every valid plan $\pi$, the V-extended state plan is valid.
A domain is trivially execution-consistent for the empty verifier $V(\cdot)=\emptyset$, as the extended state plan is equal to the original plan. A domain is also execution-consistent if the range of $V$ applied to each motion subsequence $\mu_{i}$ corresponding to action $a_{i}$ is limited to facts in $\text{Eff}(a_{i})$. In other cases, a domain can still be execution-consistent for a verifier that would introduce new facts if the domain is carefully crafted. In defining a planning domain for any task, we seek to have it execution-consistent with respect to any defined verifiers.
In our example, we want to prevent the agent from entering places that it should not, and so would like to show that the Inspection domain is execution-consistent with respect to the verifier $V_{place}$. Recall that $V_{place}$ can only introduce new VisitedPlace facts. As VisitedPlace does not appear in any action preconditions, the only way for a VisitedPlace fact to render a valid state plan invalid is to conflict with the goal specification. Consider the set of places that must be avoided to satisfy some goal state: $\mathcal{P}_{avoid}=\{p~{}|~{}\texttt{(not (VisitedPlace p))}\in\mathcal{G}\}$. If a place in $\mathcal{P}_{avoid}$ can only be visited by an action that explicitly lists it in the action effects, then the domain will be execution-consistent with respect to $V_{place}$. This can easily be guaranteed by preventing the motion planner from generating plans that enter places $\mathcal{P}_{avoid}\setminus\mathcal{P}_{param}$, where $\mathcal{P}_{param}$ is the set of places that appear as parameters to the action. In other words, we define a motion planner that cannot enter any Place which we might want to avoid, unless that Place is given as the parameter to the moveRelaxed action. As a result, we can be sure to never visit a potentially illegal place without (VisitedPlace p) being an effect of the action.
This notion of avoiding potentially illegal places implies something important with respect to the motion planning problem. If we identify a place to be redundant, and remove it from the planning domain, the definition of redundancy requires that any plan that would have passed through this place if we hadn’t removed it can be replicated. As a result, while the discrete PDDLStream planner does not need to be aware of pruned places, the motion planner we implement to solve for trajectories must still be able to plan paths through these places. Our motion planner must be aware of the full scene graph, regardless of what is pruned from the PDDLStream problem instance. We address this further in Sec. III-C. With that in mind, we can now consider which specific places we can remove from the PDDLStream problem instance.
Proposition 2 (Redundant Places).
Consider a problem instance in the Inspection Domain with no quantifiers that can be parameterized by a place in the goal. A place $p$ is redundant if no facts parameterized by $p$ appear in the initial or goal states, or if (not (VisitedPlace $p$)) appears as a clause in the conjunctive normal form (CNF) of the goal specification.
Proof.
First note that no actions in this domain have universal quantifiers, so we only need to check Definition 1.1 to show that a symbol is redundant. Consider a place $p$ such that (not (VisitedPlace $p$)) appears in the CNF of the goal. If $p$ parameterizes moveRelaxed, then (VisitedPlace $p$) is in the effects, violating the goal. Since moveRelaxed is the only action that can be parameterized by a place, no plan can parameterize $p$ and Definition 1.1 is trivially satisfied.
Next, consider a place $p$ that does not parameterize any initial or goal facts. For any plan $\pi$ with an action parameterized by $p$, let $a_{k}^{\mathcal{P}}$ denote an action parameterized by places $\mathcal{P}$, including $p$. Plan $\pi^{\prime}$ where $a_{k}^{\mathcal{P}}$ is replaced by $a_{k}^{\mathcal{P}\setminus p}$ is also valid, since state plans $\mathcal{I}_{\pi}$ and $\mathcal{I}_{\pi^{\prime}}$ only differ by a (VisitedPlace $p$) fact, and no action preconditions or goals involve this fact. Since $p$ was not in $\mathcal{P}_{avoid}$, the command sub-sequence corresponding to $a_{k}^{\mathcal{P}}$ is also valid for $a_{k}^{\mathcal{P}\setminus p}$. Thus, for any $\pi$ parameterized by $p$, we can construct $\pi^{\prime}$ that has an equivalent motion sequence but does not parameterize $p$, showing $p$ is redundant. ∎
We have now identified a potentially large (depending on the sparsity and geometry of the given scene graph) set of Place symbols that are redundant to our instantiation of the Inspection domain. Removing these places from our task planner’s domain, assuming the motion planner is still aware of them, should enable our solver to find valid plans that are guaranteed to have also been valid in the un-pruned problem. Moreover, the ability to prune these elements does not restrict the type of goals we are able to specify to our agent, preserving expressivity, while enabling planning at a larger scale. Our explicit method of defining our problem’s initial state is as follows:
Remark 1 (Problem Initialization).
In light of Proposition 2, we only include the following places when instantiating a problem: 1) the initial place that the robot is in and 2) any place that appears in the goal. A place $p$ that parameterizes a negated fact (not (VisitedPlace $(\texttt{not }(\texttt{VisitedPlace }p))$ that appears as a clause in the CNF of the goal specification can also be removed. Note that if (not (VisitedPlace $(\texttt{not }(\texttt{VisitedPlace }p))$ appears in the goal, but not as a standalone clause in the CNF, we do include that place. Any problem in the Inspection domain with this instantiation is guaranteed to return only plans that are execution-consistent with respect to the $V_{place}$ verifier discussed earlier, and as such, no plans need to be verified at runtime.
In implementing the resulting domain, there is a tradeoff between branching factor and required plan depth. Letting each move action be parameterized by $N$ places results in a search problem with large branching factor and expensive task planner preprocessing when running the search. In exchange, a single action can move through several places that appear in $\mathcal{P}_{avoid}$. Alternatively, we can restrict each move action to include only two places – those at the beginning and end points of the action – and preserve the set of solution trajectories. This comes at the expense of more numerous sub-problems and longer action plans, though we find this is a worthwhile trade off in the domains we consider. We can mitigate this computational cost by using an intermediate abstract planner to accelerate motion planning, which we describe below.
III-C Motion Planning Through Places
Critical to the success of our approach is the implementation of a motion planner that is capable of planning over large distances, while avoiding specific regions as specified by the task planner. Solving for a valid plan for a TAMP problem in a large environment requires solving for trajectories between potentially many different possible start and end poses, so the efficiency of the motion planner has a large effect on the overall runtime of the planner. The connectivity information between places in the scene graph is a useful tool for focusing the search space for feasible trajectories over large distances, since each place represents a region of navigable space. However, the connectivity information in the scene graph does not account for robot kinematic constraints or necessary standoff distances from obstacles. Thus, though we leverage information from the scene graph when possible, we still need to run a kinematic motion planner.
To find a motion plan between two configurations $c_{1}$ and $c_{2}$, we plan through the places layer of the complete, unpruned scene-graph, using A* to find a sequence of places $[p_{1},...p_{N}]$ with corresponding centers $[x_{1},...x_{N}]$ such that $c_{1}$ is in place $p_{1}$, $c_{2}$ is in $p_{N}$, and edge $(p_{i},p_{i+i})$ is present in the scene graph. Though we have pruned the PDDL planning domain to accelerate the discrete component of search, we still retain the full information in the scene graph in order to generate trajectories. This is a much simpler search problem than using the PDDL domain of the full scene graph. Second, a kinematic planner plans a path through these positions. This path ignores obstacles that may be along the path and does not necessarily stay within places on the reference path, so it is very fast to generate. For a robot with a kinematic bicycle model, this initial path might be a series of Dubins curves calculated between sequential place centers. Next, this path is checked for collisions against the geometry of all objects in the scene. Any segments of the trajectory that are rendered infeasible by collision with objects or overlap non-traversable places are re-solved by a planner that considers the obstacles, such as RRT [14]. This two-phase approach allows for fast solving of “easy” portions of the trajectory, while being able to spend time searching for feasible solutions to more difficult parts of the path. Better alignment between which edges are present in the scene graph and kinematic feasibility for the robot leads to better performance of this heuristic.
This coupling of the motion planner with the higher-level discrete planner using the places layer of the scene graph, gives us in actuality a three level hierarchical planner: 1) an abstract discrete task planner for interacting with the environment using the PDDL domain based on the pruned scene graph described in Sec. III-A, 2) a navigation planner for moving from place to place in the places layer of the full scene graph, and 3) a low-level motion planner using the abstract plan and the geometry of the scene graph to produce kinematically feasible trajectories. This planner architecture aligns naturally with the scene graph hierarchy, and we notice that with this factorization, each planning problem we care about becomes considerably simpler. The top-level planner only needs the subset of relevant the places layer, so we get a speedup by eliminating the parts of the scene graph that are redundant and reducing the planning horizon. The abstract navigation planner only plans through the places layer to one sub-goal at a time, and so can take advantage of Euclidean distance heuristics. Finally, the fine-grained motion planner is guided by the solution of the plan through the places layer, and so does not suffer from the usual challenges of finding motion plans over large distances. Crucially, we are able to do this without restricting the types of goals we can specify to the planner.
III-D Weakly Redundant Symbols and Object Pruning
We have demonstrated the ability to identify elements of a planning instance that are provably redundant when searching for a plan. However, there may be cases where planning could be made vastly more efficient by pruning elements of a problem that do not meet the strong condition of Definition 1. Consider, for example, the case where a robot is tasked with travelling across a large scene graph in the Inspection domain, containing many suspicious objects. The robot can inspect and neutralize an obstacle to pass it safely, but depending on the geometry of the scene, there may exist a path to the goal that is completely free of any obstacles. Alternatively some subset of the objects may be needed to be cleared to find safe passage. Regardless of that reality, none of the objects in the scene meet our definition of redundancy due to the fact that motion sequences which involve inspecting an object cannot be generated once the object is removed by the domain. The conditions for a symbol being redundant are quite strong – every trajectory must be preserved when removing the symbol. We now introduce a weaker notion of redundancy.
Definition 4 (Weakly Redundant Symbols).
A symbol x is weakly redundant if both of the following hold:
  1. 1.
    For every plan $\pi$ where x parameterizes an action, there is another plan $\pi^{\prime}$ that is valid where x is not an action parameter.
  2. 2.
    No action preconditions or goals, expressed in negative normal form, contains a universal quantifier that can be parameterized by x.
In contrast to strongly redundant symbols, removing a weakly redundant symbol from a planning instance may limit the set of motion sequences that can be found. An object is weakly redundant if a valid plan can be found without that object being necessary to the plan (e.g., if the robot can find a path around the object in question). Because the applicability of this definition may depend on the geometry of the scene, it is more difficult to identify symbols with weak redundancy, though it potentially applies to a wider set of objects. Consider two sets of symbols, $\mathcal{O}_{1}\subseteq\mathcal{O}_{2}\subseteq\mathcal{O}$. If there is a valid plan $p$ that is only parameterized by objects in $O_{1}$, then the objects in $O_{2}\setminus O_{1}$ are weakly redundant and do not need to appear in problem instance in order to find a valid plan.
Unfortunately, identifying weakly redundant objects in the Inspection domain can be as hard as solving the original planning problem. However, we can still take advantage of the sparsity that exists in many scene-graphs by considering an incremental approach for identifying a subset of objects without (some) weakly redundant objects. We begin planning by including some subset of all objects, and attempt to solve the planning problem. If this limited problem has a valid solution which is also a valid solution to the original problem, then we have found a plan (and have shown the excluded objects were in fact weakly redundant). Otherwise, we incrementally add objects to the planning problem, and repeat (Algorithm 1).
Identifying the initial set of objects $\mathcal{O}_{S}$ is not trivial. These objects are those that we are certain are neither redundant, nor weakly redundant. We have proven in the Inspection domain to be able to identify Places from the scene graph which are not redundant as described in Remark 1, so those are added to the problem. Any Object in the scene graph is potentially weakly redundant, depending on scene-geometry, unless it is explicitly mentioned in the goal, so we only add those goal Objects to the initial problem instance. For general problems, there is a body of literature dedicated to identifying object relevance, either by a reachability analysis [4, 5], or otherwise learning to predict importance [19] which could also be used to augment this initial set.
We then initialize a problem instance with these objects $\mathcal{O}_{I}=\mathcal{O}_{S}$. Let $\mathcal{O}_{R}$ denote the rest of the objects in the scene graph. Next, in accordance with the adaptive algorithm in Garrett et al. [6], we optimistically assume that certain sub-problems are solvable, and identify candidate action-plans which, if this assumption were true, would be a valid plans. We refer to these candidates as a plan skeletons. The set of plan skeletons for the problem instance involving objects $\mathcal{O}_{I}$ are generated by GetSkeletons (Line 7). SolveSubProblems (Line 8) attempts to solve the relevant sub-problems in a given candidate plan skeleton. If a sub-problem fails to be solved, we query the solver for feedback information (which objects in the full set $\mathcal{O}_{R}$ are a potential cause of failure). If all sub-problems were solved successfully, we return the full valid plan. Otherwise, we choose to add one or more new objects to the problem with GetNewObjects, which considers the feedback about the failure mode that CheckSolution (Line 9) returned. This approach will continue to incrementally add objects until a plan is found, and so is complete [^1] Assuming SolveSubProblems is complete and GetSkeletons will return arbitrarily long skeletons once all objects have been added..
Algorithm 1 Incremental Object Solver
1:procedure IncrementalObjectSolver($A,S,I,G$)
2:     $\mathcal{O}_{S}\leftarrow GetRelevantObjects(\mathcal{O})$triangleright Find objects relevant for non-collision reasons
3:     $\mathcal{O}_{I}\leftarrow\mathcal{O}_{S}$
4:     $\mathcal{O}_{R}\leftarrow\mathcal{O}\setminus\mathcal{O}_{S}$
5:     while $|\mathcal{O}_{I}|<|\mathcal{O}|$ do
6:         $SkeletonInfo=[~{}]$
7:         for $k\in GetSkeletons(A,S,\mathcal{O}_{I},G)$ do
8:              $T\leftarrow SolveSubProblems(k)$
9:              $Feedback\leftarrow CheckSolution(T,\mathcal{O}_{R})$
10:              $SkeletonInfo.\texttt{append}(Feedback)$
11:         end for
12:         if $\pi\in SkeletonInfo$ is valid then
13:              return $\pi$
14:         else
15:              $\mathcal{O}_{new}=GetNewObjects(C,F)$
16:              $\mathcal{O}_{I}=\mathcal{O}_{I}\cup\mathcal{O}_{new}$
17:              $\mathcal{O}_{R}=\mathcal{O}_{R}\setminus\mathcal{O}_{new}$
18:
19:         end if
20:     end while
21:     return INFEASIBLE
22:end procedure
Existing literature related to identifying irrelevant objects in a planning problem often focuses on ignoring or more efficiently encoding symbols based on careful analysis of the PDDL problem’s logical factorization  [10, 5]. In contrast, our algorithm is well suited to address the geometric dependence of movement actions on objects that may render a trajectory infeasible. A movement action must avoid collision with all objects in the environment, and it is difficult to show that an object cannot interfere with any plan and can be ignored without explicitly inspecting the geometry. Algorithm 1 can be seen as a way of lazily uncovering which of these objects do interfere with plans that would otherwise be valid.
We take advantage of the hierarchy in the scene graph to implement GetNewObjects, which is tasked with identifying which new objects to add to the planning problem based off feedback from failed sub-problems. Each time we attempt to solve a motion plan, we produce a plan that avoids the objects that have been added to our planning domain (if one exists). Then, we check that plan against the full set of objects. If one of those were intersected, and we cannot find a plan that does not intersect any of the full set of objects, the first object is potentially relevant to the planning problem, and we add it to the planning domain. By adding this object to the planning problem, the planner can now consider inspecting and neutralizing it, thus introducing a new potential path to a goal state. We envision many potential extensions to this approach, such as learning parameters for these subroutines, which we do not explore in this work. In the next section we evaluate our planning approach in the Inspection domain.
IV Evaluation
Figure 2: Three of the maps used for evaluation (not including the KITTI environment). Top Left: narrow alley map. Top Right: A simple 10x10 grid world map. Bottom: Scene graph built from real data collected by a robot in an office environment.
There are three primary axes of complexity for planning problems in large scene graphs that we are interested in investigating. One is the complexity of the goal specification, another is correlated to the scale of the environment, and the third is related to the geometry of the scene (and potential obstructions therein). In our experiments, we aim to identify which types of planning problems our planning approach is well suited for considering these sources of complexity. To that end, we compare our encoding of the Inspection domain to the dense, direct encoding in a variety of different settings. We test on four map archetypes (Fig. 2)– a synthetic small constrained alleyway, a synthetic 10x10 gridworld, a scene graph built from real data in an office environment comprising 557 Places and 28 Objects, and a much larger scene graph built from the KITTI dataset composed of 17861 Places and 1315 Objects (Figs. 1 and 2). For each environment, we test several different goal clauses across different variations in robot and object initial conditions.
To randomize tasks across trials, we define a mechanism for sampling goal specifications according to an increasing number of clauses in Disjunctive Normal Form (DNF). More explicitly, we generate goals in DNF with N clauses, where each clause has K conjunctions. For example, if N = 2 and K = 3, the goal looks like this: (Or ($C_{1}$, $C_{2}$)), where we sample 3 clauses for $C_{i}$, e.g., (And ((Visited P1), (Safe O4), (Not (Visited P9)))).
IV-A Scene Graph Size
Figure 3: This log-log plot compares the time (in seconds) to solve tasks of comparable complexity across different environmental scales for the dense formulation and the proposed sparse formulation. In the smallest environment, the dense baseline planner performs just as well as our approach. However, as we increase the scale, first to a 10x10 grid, and then to a real scene graph constructed on the floor of an academic building, we see the benifits of our approach.
First, we investigate the effect of scene graph size on our ability to plan to a set of goals of consistent complexity. For this set of trials, our goal complexity is (N, K) = (3, 3), and we compare planning time for the direct encoding to our planner, highlighted in Fig. 3. Each point on the scatter plot corresponds to a single trial and, as highlighted in the legend, different colors correspond to different environment types. We begin in the Alley environment which, as shown in Fig. 2, is a 2x6 grid world containing only connected places. In this environment, for this level of complexity, our planner performs about as well as the dense encoding, as there is not much advantage to sparsification in such a small environment. As we scale up however, this changes. The next environment we consider is a 10x10 grid, where we see modest improvement as shown in Fig. 3. In the figure, any samples above the black line indicate that our planner outperforms the dense baseline. As we scale up even further, into the indoor DSG, we see the baseline planner taking on the order of hundreds of seconds to plan on average, while our planner averages in the tens of seconds. We note that we only considered goals that involved visiting or not visiting certain Places in the graph. When we attempted to introduce object inspection, the dense planner timed out before finding a plan in all instances. Similarly, when testing the baseline in the KITTI scene graphs, it was also unable to find solutions for goals of any complexity. Planning with these relatively simple goals, our planner experienced only a modest increase in planning time as the size of the map scaled.
IV-B Goal Complexity
Figure 4: Here we highlight how our approach scales with the complexity of the goal specification in the simple 10x10 grid world. As we increase the number of unique PDDL objects in the goal specification, the problem is no longer sparse, and so no longer benefits from our approach. Eventually, the dense, direct encoding begins to outperform our method.
Next, we consider the effect of increasing goal complexity on planning time. To do this, we investigate a series of various different goal constructions in the Grid World environment. Specifically, we run experiments for K = 5 and K = 10, incrementing the N value from 1 to 5 for each K value. Stated differently, we have conjunctive clauses that are either 5 or 10 atoms long, and “Or” together 1, 2, 3, 4, or 5 of these conjunctions. Each of these clauses is once again either (Visited P1) or (Not (Visited P1)). Figure 4 presents a plot comparing the complexity of the goal in terms of total unique symbol referenced vs planning time. For less complex goals in this environment, our planner outperforms the dense planner, up to a crossover point at around 20 unique objects. At this point, the benefits of sparsity disappear for a scene of this size. Notably however, if we were to introduce obstructing objects, the dense direct encoding baseline is unable to solve goals of any meaningful complexity. Once again, not included in this plot are the results on the KITTI dataset, as the direct encoding never successfully completes a trial in this setting due to timing out.
IV-C Object Obstruction in KITTI
Figure 5: An example plan from the KITTI environment. The robot begins in the bottom left corner and is tasked with inspecting one object (denoted by the red triangle at the end of the black trajectory). Along the way, there are numerous objects potentially blocking the agent’s path, so it must add at least one to its planning domain. After inspect and neutralizing this object, the robot is free to reach its goal.
Finally, we investigate the performance of the incremental object solver algorithm in task instances where objects not directly listed in the goal must be inspected in order to solve the task. In this experiment, we give a robot one of two goal types in a scene graph built from the KITTI dataset (Figs. 1 and 5): either (Visited $P_{i}$) or (Safe $O_{j}$). Given the size of the map, satisfying these goals may require the agent to traverse a large distance, but more importantly, if there are obstructing objects in the way, it may be forced to inspect and neutralize them to find a safe path to the goal. As a baseline, we sample 20 goals in the map shown in Fig. 5 using our planner without any of the objects being labeled as suspicious. In this case, they do not obstruct the agent’s path, and we find plans in 19 of 20 trials.
Next, we “activate” 13 objects in the scene by labeling them as suspicious. A suspicious object has an inflated radius that is only safe for the robot to enter after it has been inspected and neutralized (in the KITTI scene, this radius is large enough to block an entire road as shown by the blue objects in Fig 5). For the agent to inspect the object, it has to sample a pose that is traversable and within line-of-sight and range of the object. Then, by taking the inspect action, the object is now safe, and can be passed. To demonstrate the difficulty of this task, and to highlight the importance of object pruning, we attempt to solve these same tasks without using our incremental feedback approach for pruning potentially weakly redundant objects (Sec. III-D). Instead, we add all suspicious objects to the scene directly. Using this encoding, the planner only succeeds in finding a plan in 4 out of 20 trials. Inspecting these solutions further reveals that in all 4 of these successful cases, there was a direct path to the goal without inspecting any objects. This result makes sense, as the odds of sampling the correct object to inspect is low without the benefit of geometric information.
Finally, we test our proposed approach with pruning objects from the planning domain, and incrementally adding them back if they impacted the failure of a sub-problem during search (Sec. III-D). Our planner is able to solve 12 of the 20 trials, including 9 cases wherein the agent inspected one or more obstructing objects on the way to its goal. These experiments further demonstrate the importance of our proposed approach to sparsifying otherwise dense, long-horizon planning problems. An example plan, where the agent investigates two objects on the way to its goal is shown in Fig. 5.
V Related Work
There has substantial recent work enabling construction of information-rich scene graphs. Wu et al. [23] use a graph neural network to build higher-level abstractions from sensor data. Hughes et al. [12] and Bavle et al. [3] build scene graphs while localizing a robot’s motion, allowing for co-optimization of pose and scene estimation. Other recent directions include the registration of scene graph pairs [17].
There has been recent work in the area of deriving PDDL representations for task planning from scene graphs [1], with a particular focus on using the hierarchical nature of scene graphs to sparsify the representation in order to make planning tractable. The approach used by Agia et al. [1] is only guaranteed to produce valid solutions for very specific planning domains, where only constraints between symbols with a clear “ancestor” relationships are expressible, and it is unclear how to extend this to a more general set of planning tasks. Unfortunately the scene graph abstraction may not satisfy the property of downward-refinement [2], breaking many of the assumptions in task planning. The existance of low-level geometric constraints requires going beyond task planning approaches, and into the realm of TAMP.
There has been additional work in pruning superfluous elements of scenes to accelrate TAMP. Silver et al. [19] learn to predict which symbols are relevant to a particular TAMP problem. Similarly, Srivastava et al. [20] attempt to guide TAMP from failed motion plans, much like we do in Sec III-D, however do so by adding additional goal conditions, which can complicate planning.
VI CONCLUSIONS
In this work we proposed an approach for enabling and accelerating TAMP in large scene graphs. To do this, we defined characteristics of planning domains that permit the pruning of certain symbols. Then, we proposed a method for deriving a domain from a Hydra scene graph which has these characteristics, and demonstrate how we prune Places and Objects from the domain. We also prove how the plans we produce from this pruned scene graph are valid and conform to the constraints of the full planning domain. Finally, we demonstrate experimentally how our approach scales with scene graph size, goal complexity, and geometric constraints in several environments, including a scene graph built from the KITTI dataset.
In future work, we hope to demonstrate under what conditions we can extend our pruning method to other domains derived from large-scale scene graphs. Furthermore, augmenting our approach with learned methods for object pruning is a natural extension. The metric-semantic information in the scene graph is potentially a strong signal for a learner to identify further irrelevant symbols in a planning domain.
Acknowledgements
This work was partially funded by the ARL DCIST program, by Lincoln Laboratory’s Autonomy Al Fresco program, and by the National Defense Science and Engineering Graduate Fellowship program.